OpenStudy (lopus):

I have tried to create tables with latex and I've failed and it happens in forums like here

OpenStudy (lopus):

\begin{tabular}{ l c r } 1 & 2 & 3 \\ 4 & 5 & 6 \\ 7 & 8 & 9 \\ \end{tabular}

OpenStudy (lopus):

\begin{tabular}{ |l|l| } \hline \multicolumn{2}{|c|}{Team sheet} \\ \hline GK & Paul Robinson \\ LB & Lucus Radebe \\ DC & Michael Duberry \\ DC & Dominic Matteo \\ RB & Dider Domi \\ MC & David Batty \\ MC & Eirik Bakke \\ MC & Jody Morris \\ FW & Jamie McMaster \\ ST & Alan Smith \\ ST & Mark Viduka \\ \hline \end{tabular}

OpenStudy (kinggeorge):

\[\begin{array}{ l c r } 1 & 2 & 3 \\ 4 & 5 & 6 \\ 7 & 8 & 9 \\ \end{array}\]I would suggest using "array" instead of "tabular."

OpenStudy (e.mccormick):

The MathJax is limited in what modules it supports. The good news is, many of the formatting tools will work in array for changing justification and adding lines. \[\begin{array}{|rc|l|} \hline1 & 820 & 3\\ \hline740 & 5 & 960\\ \hline \end{array}\] Also, if you really need a little more range of things, you could use this: http://www.codecogs.com/latex/eqneditor.php?lang=en-en Then you can link it: http://latex.codecogs.com/gif.latex?%5Cdpi%7B120%7D%20%5Cbg_white%20%5Cbegin%7Btabular%7D%7B%20%7Cl%7Cl%7C%20%7D%20%5Chline%20%5Cmulticolumn%7B2%7D%7B%7Cc%7C%7D%7BTeam%20sheet%7D%20%5C%5C%20%5Chline%20GK%20%26%20Paul%20Robinson%20%5C%5C%20LB%20%26%20Lucus%20Radebe%20%5C%5C%20DC%20%26%20Michael%20Duberry%20%5C%5C%20DC%20%26%20Dominic%20Matteo%20%5C%5C%20RB%20%26%20Dider%20Domi%20%5C%5C%20MC%20%26%20David%20Batty%20%5C%5C%20MC%20%26%20Eirik%20Bakke%20%5C%5C%20MC%20%26%20Jody%20Morris%20%5C%5C%20FW%20%26%20Jamie%20McMaster%20%5C%5C%20ST%20%26%20Alan%20Smith%20%5C%5C%20ST%20%26%20Mark%20Viduka%20%5C%5C%20%5Chline%20%5Cend%7Btabular%7D Which may be ugly at times, so you could also download it as a gif/png (change background to white!) and then save than and attach it.

OpenStudy (lopus):

what mean {l r c}

OpenStudy (e.mccormick):

Left Right Center

OpenStudy (lopus):

thank you

OpenStudy (ybarrap):

$$\begin{array}{|rl|c|} \hline ^a_b&_b^{\Huge{a}}&c\\ \hline \end{array}$$